The assistant manager of a surf shop estimates that 65% of customers will make a purchase. Part A: How many customers should a salesperson expect until he finds a customer that makes a purchase?
Answer #1

The assistant manager of a surf shop estimates that 65% of customers will make a purchase

expected number of customer = 1/p = 1/0.65 = 1.54